OpenStudy (anonymous):

A company's monthly profit, P, from a product is given by P = -x2 + 105x - 1050, where x is the price of the product in dollars. What is the lowest price of the product, in dollars, that gives a monthly profit of $1,550? (Do not put a $ in your answer.)

OpenStudy (anonymous):

This is one of those problems where you just have to look at what you have and what you are solving for. If you have a profit of 1550, then that means you have a value of P in the equation of 1550 like so: \[1550=-x^2+105x-1050\]

OpenStudy (anonymous):

Now, you need to put the quadratic equation in a form that you can factor and solve, or you may use any number of methods to solve it (I like the quadratic formula), but lets see if we can factor it.

OpenStudy (anonymous):

You always want it to be in the form 0=\[ax^2+bx+c\]

OpenStudy (anonymous):

To do that, simply subtract 1550 from both sides.

OpenStudy (anonymous):

\[-x^2+105x-2600\]=0

OpenStudy (anonymous):

Multiply both sides by (-1) to make the leading term positive (just to make the math a bit easier in the end when we solve this thing), \[x^2-105x+2600=0\]

OpenStudy (anonymous):

Give me a second to see if I can factor this.

OpenStudy (anonymous):

I dont think it would be easy to do, I would solve with quadratic formula.

OpenStudy (anonymous):

So, using the quadratic formula, we know that a = 1, b = -105 and c = 2600

OpenStudy (anonymous):

The quadratic formula becomes: \[(-(-105)\pm \sqrt{(-105)^2-4(1)(2600)})/2(1)\]

OpenStudy (anonymous):

= x, and the problem asks you to choose the smallest value of x

OpenStudy (anonymous):

so the two solutions are x = \[(105+\sqrt{(-105)^2-(4)(1)(2600)} )/2\]

OpenStudy (anonymous):

And \[(105-\sqrt{(-105)^2-(4)(1)(2600)} )/2\]

OpenStudy (anonymous):

I gotta figure out how to type up fractions with this thing

OpenStudy (anonymous):

\[\frac{(105+\sqrt{(-105)^2-(4)(1)(2600)}}{2}\]

OpenStudy (anonymous):

And \[\frac{(105-\sqrt{(-105)^2-(4)(1)(2600)} )}{2}\]

OpenStudy (anonymous):

Pick the smallest answer, and see if it works. Does this make sense?

OpenStudy (anonymous):

If it does not work, I may have messed the arithmetic up

OpenStudy (anonymous):

to me it dont sense

OpenStudy (anonymous):

Are you familiar with the quadratic formula?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Oh, well I can give you a quick intro on it if you would like.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So what it is, is it is a way of solving equations that are in the following form: \[ax^2+bx+c=0\]

OpenStudy (anonymous):

An example of an equation in this form is something like: \[2x^2+3x-9=0\]

OpenStudy (anonymous):

So let me ask you a quick question to make sure you are up to speed. If I have an equation \[3x^2-5x+6=0\]

OpenStudy (anonymous):

What are the values of a, b, and c?

OpenStudy (anonymous):

since it is in the form \[ax^2+bx+c=0\]

OpenStudy (anonymous):

a=3 b=5 c=6

OpenStudy (anonymous):

Very close :), in this example you have to be careful, because the value of b is actually -5, not positive 5.

OpenStudy (anonymous):

right, because it is really \[3x^2+(-5)x+6=0\]

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

The other two are correct.

OpenStudy (anonymous):

Cool, so the quadratic equation is just a big fancy formula that was figured out. And if you know the values of a, b, and c, you can just plug them in and you will get the solutions of x.

OpenStudy (anonymous):

It is in this form:

OpenStudy (anonymous):

x=\[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Have you seen this before in your class?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok cool, lets do an easy example of how you plug in the values to this thing. So here is another equation in the form \[ax^2+bx+c=0\]

OpenStudy (anonymous):

You tell me what a,b, and c are. \[2x^2+4x-1=0\]

OpenStudy (anonymous):

a=2b=4c=-1

OpenStudy (anonymous):

Yup, good. Now here is how you plug them into the formula.

OpenStudy (anonymous):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}=\frac{-4\pm \sqrt{(4)^2-4(2)(-1)}}{2(2)}\]

OpenStudy (anonymous):

Look at it very closely, where there was a "b", I just plugged in the value that we got for b in the equation, and the same for the values of a and c

OpenStudy (anonymous):

Once you understand that, I will tell you how to deal with the

OpenStudy (anonymous):

plus or minus sign

OpenStudy (anonymous):

i get that

OpenStudy (anonymous):

Ok, most people get tripped up by the plus or minus sign, and all it means is that there are two ways to write it. You can write the solution with the plus sign, and you can write the solution with the minus sign. So the one that we just found, to get rid of the plus or minus sign, just split it up like so: \[x=\frac{-4+\sqrt{(4)^2-4(2)(1)}}{2(2)}\]

OpenStudy (anonymous):

And \[x=\frac{-4-\sqrt{(4)^2-4(2)(1)}}{2(2)}\]

OpenStudy (anonymous):

Thats all there is to it, I recommend always when you substitute the values of a, b, and c into the equation that you put them in parenthesis because remember that for example \[-2^2\neq(-2)^2\]

OpenStudy (anonymous):

Now you would just have to simplify the answers to get x

OpenStudy (anonymous):

okay now i get it thank you

OpenStudy (anonymous):

\[x=\frac{-4+\sqrt{8}}{4}=-1+\frac{\sqrt{8}}{4}\]

OpenStudy (anonymous):

and the same for the other solution

OpenStudy (anonymous):

Awesome, I hope you do. That is all you need to do for your problem, the only trick is getting it in the right form.

OpenStudy (anonymous):

Try your problem now, and let me know if you have any problems